# Back from the dead: AT2019aalc as a candidate repeating TDE in an AGN

### v1

@@ Abstract
Methods. We conducted a multi-wavelength monitoring program (from radio to X-rays) of AT2019aalc during its re-brightening in 2023.
@@
Tidal disruption events (TDEs) are rare transient events, which occur when a star approaches a supermassive black hole (SMBH) and the tidal forces of the latter rip the star apart (Rees 1988). About half of the star’s material is accreted around the black hole, generating a luminous outburst across the electromagnetic spectrum. Despite most of these transient sources have been discovered by optical telescopes, the origin of their optical/UV emission is poorly understood. One model is colliding stellar debris stream shocks (Piran et al. 2015). Another scenario is reprocessed soft X-ray/far UV emission (originating from the debris accretion onto the black hole) by an extended elliptical disk, or an envelope, or an outflow surrounding the central debris disk (for a recent review see Bu et al. 2022). TDE optical light curves are characterized by quick and colorless evolving but also fast (on monthly timescales) decaying optical emission. Several TDEs were detected in X-rays with a very soft X-ray spectrum typically explained by late-time accretion disk formation (Saxton et al. 2020). In addition, a handful of TDEs produced detectable radio emission, explained by delayed111Note that the SMBH accretion rate does not peak on the same timescale as the mass fall-back rate. non-relativistic outflows (Horesh et al. 2021; Cendes et al. 2023), or only in instances by newly launched relativistic on-axis jets (Berger et al. 2012; Brown et al. 2015; Cenko et al. 2012; De Colle & Lu 2020; Yao et al. 2023a). Spectroscopically, TDEs are characterized by a strong blue continuum and broad emission lines ( km s-1) e.g., H and H. As the photosphere shrinks over time, the flux of the ionized lines such as He ii tend to increase while the He ii/H ratio increases (e.g., Hung et al. 2017; Charalampopoulos et al. 2022a) and, in a few cases, Bowen Fluorescence (BF; Bowen 1934, 1935) lines e.g., N iii at appear (e.g., Blagorodnova et al. 2019; Charalampopoulos et al. 2022a).
To date, roughly promising candidate TDEs are known. However, since most of the above-mentioned observed properties do not exclusively characterize TDEs, spectroscopic coverage is required to distinguish TDEs from ‘TDE impostors’ such as conventionally accreting active galactic nuclei (AGN), changing-look AGN (CLAGN), supernovae (SNe) or even stellar collisions (Zabludoff et al. 2021).
@@
TDEs generally result in one optical flare, however, in a few cases a re-brightening episode (or episodes) was detected even years after the initial flare. These might be explained in the following way. Under special circumstances, a star on a grazing orbit might be partially disrupted and only a fraction of the stellar mass will be deposited onto the SMBH. The secondary disruption of a surviving core after it reaches the pericenter again results in X-ray re-brightening explained by renewed accretion. The first systematically identified repeating partial TDE, AT2020vdq, has been classified after simultaneous multi-wavelength monitoring (Somalwar et al. 2023). Only a few other partial TDEs or candidates are known to date, these are ASASSN 14ko (Payne et al. 2021); eRASSt J045650.3-203750 (Liu et al. 2023); RX J133157.6-324319.7 (Hampel et al. 2022); AT2018fyk (Wevers et al. 2019a, 2023); AT2019avd (Chen et al. 2022) and AT2022dbl (Lin et al. 2024).
@@
Based on the observed properties of TDE host galaxies, French et al. (2020) estimated a TDE rate of per year per galaxy which is well borne out by recent data (Yao et al. 2023b). Importantly, a similar fraction of TDEs should take place in AGN hosts (e.g., Chan et al. 2019; Ryu et al. 2024). The multi-wavelength characteristics of such a TDE should clearly differ from the ones taken place in quiescent galaxies, as the AGN is surrounded by a pre-existing accretion disk. Due to the extreme difference in density between the two, neither the star nor the accretion disk are affected after the first passage through the disk. However, the confined debris stream has such a low density after the star is tidally disturbed that when it interacts with the disk the collision between the stream and the disk can potentially dissipate much of the kinetic energy possessed by the stream and the disk gas near the impact point, and it can be radiated away (Chan et al. 2019). These transients are challenging to distinguish from AGN variability. Only a few candidate TDEs taken place in AGN (in the following: TDE-AGN) were studied earlier (e.g., PS16dtm - a TDE in a Narrow-Line Seyfert-1 (NLSy1) galaxy (Blanchard et al. 2017) or the multiple soft X-ray flares in IC 3599 (Campana et al. 2015) and GSN 069 (Shu et al. 2018)). In contrast with TDEs seen in earlier quiescent galaxies, these transients are characterized by slowly decaying optical emission and bumps tend to appear after the peak.
@@
The family tree of AGN-related transients became much broader in the past years (e.g., Frederick et al. 2021). In addition to TDEs, flaring NLSy1s, CLAGN and different types of SNe were classified as typical nuclear transients. Recently, Trakhtenbrot et al. (2019) identified a new class of flares from accreting SMBHs. In , the transient events AT2017bgt, F01004-2237 and OGLE17aaj have been identified as the first members of the class based on their unique optical spectroscopic features. Later, the peculiar X-ray loud transient AT2019avd (Trakhtenbrot et al. 2020) was found with remarkable common properties, however, its classification as an AT2017bgt-like transient is unclear. A more promising case is AT2021loi (Makrygianni et al. 2023), a UV-bright transient event with very similar multi-wavelength properties to the instances studied by Trakhtenbrot et al. (2019). Those show unobscured AGN-like optical spectra with significant and persistent BF lines, extremely strong UV emission and slowly, on yearly scales, decaying optical emission. These AGN are characterized by significantly and longer-term intensified accretion and might be powered by a binary TDE or an interaction between an outflow and the broad-line region (BLR) or binary SMBHs (Trakhtenbrot et al. 2019). Notably, these transients take place in different types of AGN; AT2017bgt and OGLE17aaj in narrow-line AGN, AT2021loi in a broad-line AGN, F01004-2237 in an ultra-luminous infrared galaxy (ULIRG) while AT2019avd occurred in an earlier quiescent galaxy. Hereafter, we will refer to this new class of AGN flares as Bowen Fluorescence Flares (BFFs), following Makrygianni et al. (2023). Using the Zwicky Transient Facility (ZTF, Masci et al. 2019) Public Survey data, Dgany et al. (2023) found only one BFF candidate (AT2021seu) among spectroscopically classified transients, suggesting that these flares are rare. Other candidates, namely VT J1548 (Somalwar et al. 2022) and AT2022fpx, (Koljonen et al. 2024) were found recently.
@@
In recent years, two nuclear transients were found to be spatially coincident with neutrino events detected by the IceCube Neutrino Observatory based on ZTF follow-up campaigns of IceCube neutrino alerts (Stein et al. 2023). The TDE AT2019dsg is likely associated with the IceCube neutrino, IC191001A (Stein et al. 2021) and a further promising association, between the candidate TDE AT2019fdr and the IceCube neutrino IC200530A (Reusch et al. 2022) points to TDEs as a potential new class of sources of the extragalactic neutrinos. Both sources were found to have unusually luminous mid-infrared emission, and a search for similar flares (van Velzen et al. 2024) led to the identification of the transient event AT2019aalc (a.k.a. ZTF19aaejtoy) with the neutrino alert IC191119A (see Sect. 2. for explanation).
@@
We conducted multi-wavelength observations (from radio to X-rays) during the re-brightening and compare it to the initial flare associated with the high-energy neutrino, with the aim to characterize the multi-wavelength properties of this unusual transient event and to classify it.
The paper is organized as follows. We explain the discovery of the transient event and give an insight about its host galaxy in Sect. 2, we summarize our observing campaign and the data reduction in Sect. 3, we present the results of the observations in Sect. 4 and discuss them in Sect. 5 and finally give a summary and conclude in Sect. 6. Throughout the paper, we adopt a flat CDM cosmological model with parameters , , and . In this model (Wright 2006), mas angular size corresponds to pc projected linear size at the source redshift (Ahn et al. 2012). This redshift corresponds to a luminosity distance of Mpc. All magnitudes are given in AB system (Oke 1974).
@@ 2 The transient event: AT2019aalc
Studying the infrared properties of AT2019dsg and AT2019fdr have revealed a strong reverberation signal i.e. the infrared dust echo, known as reprocessed emission of optical/UV photons in the infrared by hot dust around the SMBH, at a distance of pc (Dou et al. 2016; Lu et al. 2016; van Velzen et al. 2021b, 2024). In addition to the exceptionally high infrared luminosities (due to the infrared echo) of these transients, the neutrino detection times appear to be temporally consistent with the infrared luminosity peaks (Stein et al. 2021; Reusch et al. 2022). This discovery motivated a systematic search for neutrino emission using an extended sample of black hole flares (van Velzen et al. 2024). This archival search revealed the nuclear transient AT2019aalc with a powerful optical flare and dust echo emission in spatial coincidence with the high-energy neutrino event, IC-191119A (IceCube Collaboration 2019) detected by the IceCube Observatory.
@@
The host galaxy of AT2019aalc (2MASX J15241665+0451192) is a barred spiral (SBa) galaxy hosting an active nucleus (Oh et al. 2013). Based on its Sloan Digital Sky Survey (SDSS; Abazajian et al. 2009) spectrum it was further classified as a broad-line Seyfert-1 galaxy (Trump et al. 2013; Liu et al. 2019). The galaxy was detected in archival radio sky surveys e.g., the Very Large Array Faint Images of the Radio Sky at Twenty-cm (VLA FIRST; Condon et al. 1998). Fig. 1 shows the Pan-STARRS (Chambers et al. 2016) i-band optical image of the host galaxy together with the contours of the FIRST radio data (at GHz) and the position of the X-ray counterpart detected by Swift/XRT in . The host galaxy is not present in any X-ray catalogs prior to the optical discovery of the transient in .
@@
In addition to the unusually significant infrared echo, AT2019aalc shares multi-wavelength characteristics with the two other neutrino candidate transient events: a soft X-ray spectrum and transient radio emission (van Velzen et al. 2024). Moreover, similar to the other two instances, the infrared peak of the transient event is temporarily coincident with the detection time of the likely associated neutrino event. Notably, Winter & Lunardini (2023) estimated the highest neutrino fluence for AT2019aalc, due to its high estimated SMBH mass (, van Velzen et al. 2024) and low redshift (, Ahn et al. 2012).
Since the proposed association with the high-energy neutrino event was announced roughly years after the discovery of the transient event, the source was not at the focus of attention or being directly monitored at the time when the neutrino was emitted. Thus, the lack of spectroscopic monitoring during the first optical flare hindered a clear classification of the event. Nevertheless, out of the more than AGN detected by the ZTF, less than show similarly rapid and large outbursts (Reusch et al. 2022; van Velzen et al. 2024) suggesting a very unusual AGN case. The TDE-characteristic evolution of the blackbody light curve including a large and rapid optical flux increase is compatible with a TDE-related scenario. However, due to the host galaxy classification, extreme AGN variability provides an alternative explanation for the transient event.
Interestingly, roughly years after the first flare, a significant optical re-brightening started in mid-May 2023 (Veres et al. 2023), illustrated by the long-term ZTF difference light curve shown in Fig. 2. Although double-peaked optical light curves could be a quasi-universal signature of TDEs around massive () black holes (Wevers et al. 2019b), such a long double peaked light curve with two comparable peak magnitudes would be very untypical for a regular TDE. Throughout this paper we will discuss the results of our observations taken around and after the peak of the second optical flare in the light of different TDE,- and AGN-related scenarios.
@@ 3.1.1 ZTF
ZTF forced photometry at the location of AT2019aalc was obtained following the recommendations outlined in Masci et al. (2023). We corrected for extinction with adopting a galactic extinction of (Schlafly & Finkbeiner 2011). As shown in Fig. 2, the forced photometry yields a negative flux in the period between the two flares (and also just before the first flare). This implies that during these epochs, the flux decreased below the mean flux of the reference images. For the -band these were obtained in between March and May, 2018. While the baseline used to obtain the difference flux of the -band is based on images obtained between March and September, 2018.
@@ 3.1.2 UVOT
The Neil Gehrels Swift Observatory (Arnaud 1996a) performed Target of Opportunity (ToO) observations (project codes: 18972 and 19188, Proposers: Reusch, Veres& Sniegowska) on AT2019aalc from 2023-06-23 to 2024-05-27 with a total exposure time of ks. The Ultraviolet and Optical Telescope (UVOT, Roming et al. 2005) onboard the satellite observed the target field with the following four filters in each epoch: UW2 (central wavelength, 1928 ), UVM2 (2246 ), UW1 (2600 ) and U (3465 ). We analyzed the UVOT images using the Swift UVOT tools included in the heasoft 6.30 software package. We measured the flux using uvotsource, applying a circular aperature of arcseconds (which capture nearly all the flux of the host galaxy). The background was measured using four 6 arcsecond regions in four quadrants away from the host galaxy. To estimate the UV flux before the flare, we use GALEX (Martin et al. 2005) observations obtained in 2007 and 2011. We use the same 15” circular aperture that was applied to the Swift/UVOT observations and use the gPhoton software (Million et al. 2016) to extract the flux. Over the 4 year baseline the GALEX observations, the NUV magnitudes change by about 0.3 mag. Following van Velzen et al. (2021a), we combine the mean GALEX flux with the SDSS (York et al. 2000) magnitudes and apply Prospector (Johnson & Leja 2017) to find the best-fit stellar population synthesis model (Conroy et al. 2009) that describes these data. From this model we estimate the baseline flux in the UVOT filters and this baseline is subtracted to obtain the difference flux. Finally, we corrected for galactic dust extinction for each filter.
The baseline flux in the UVM2 band is 17.2, implying a flux increase of 1.8 mag relative to the peak of the UVM2 light curve. This large flux increase implies that, at the peak of the flare, we are not sensitive to the value of the baseline flux. This is useful because the baseline of the ZTF forced photometry was obtained around 2018 (at the start of the ZTF survey), while the baseline that is used to obtain the UVOT difference flux is mainly determined by the GALEX observations that were obtained a decade earlier.

### v2

@@ Abstract
Methods. We conducted a multi-wavelength monitoring program (from radio to X-rays) of AT2019aalc during its re-brightening in 2023/2024.
@@
Tidal disruption events (TDEs) are rare transient events, which occur when a star approaches a supermassive black hole (SMBH) and the tidal forces of the latter rip the star apart (Rees 1988). About half of the star’s material is accreted around the black hole, generating a luminous outburst across the electromagnetic spectrum. TDE optical light curves are characterized by quick and colorless evolving but also fast (on monthly timescales) decaying emission. Several TDEs were detected in X-rays with a very soft spectrum typically explained by late-time accretion disk formation (Saxton et al. 2020). In addition, a handful of TDEs produced detectable radio emission, explained by delayed111Note that the SMBH accretion rate does not peak on the same timescale as the mass fall-back rate. non-relativistic outflows (Horesh et al. 2021; Cendes et al. 2023), or only in instances by newly launched relativistic on-axis jets (Berger et al. 2012; Brown et al. 2015; Cenko et al. 2012; De Colle & Lu 2020; Yao et al. 2023a). Spectroscopically, TDEs are characterized by a strong blue continuum and broad emission lines ( km s-1) e.g., H and H. In a few cases, Bowen Fluorescence (BF: Bowen 1934, 1935) lines e.g., N iii at appear (e.g., Blagorodnova et al. 2019; Charalampopoulos et al. 2022a).
@@
TDEs generally result in one optical flare, however, in a few cases a re-brightening episode was detected, even years after the initial flare. These might be explained in the following way. Under special circumstances, a star on a grazing orbit might be partially disrupted and only a fraction of the stellar mass will be deposited onto the SMBH. The secondary disruption of a surviving core after it reaches the pericenter again results in re-brightening. Only a few partial TDEs or candidates are known to date. These are AT2020vdq (Somalwar et al. 2023), ASASSN 14ko (Payne et al. 2021), eRASSt J045650.3-203750 (Liu et al. 2023), RX J133157.6-324319.7 (Hampel et al. 2022), AT2018fyk (Wevers et al. 2019, 2023), AT2019avd (Chen et al. 2022), AT2022dbl (Lin et al. 2024) and AT2021aeuk (Sun et al. 2025).
@@
Based on the observed properties of TDE host galaxies, French et al. (2020) estimated a TDE rate of per year per galaxy which is well borne out by recent data (Yao et al. 2023b). Importantly, a similar fraction of TDEs should take place in AGN hosts (e.g., Chan et al. 2019; Ryu et al. 2024). These transients are challenging to distinguish from regular AGN variability. Only a few candidate TDEs taken place in AGN (in the following: TDE-AGN) were studied earlier (e.g., PS16dtm - a TDE in a narrow-line Seyfert-1 (NLSy1) galaxy (Blanchard et al. 2017) or the multiple soft X-ray flares in IC 3599 (Campana et al. 2015) and GSN 069 (Shu et al. 2018)). In contrast with TDEs seen in earlier quiescent galaxies, these transients are characterized by slowly decaying optical emission and bumps tend to appear after the peak.
@@
The family tree of AGN-related transients became much broader in the past years (e.g., Frederick et al. 2021). In addition to TDEs, flaring NLSy1s, changing-look AGN (CLAGN) and different types of SNe were classified as typical nuclear transients. Recently, Trakhtenbrot et al. (2019) identified a new class of spectroscopically unique flares from accreting SMBHs. In , the transient events AT2017bgt, F01004-2237 and OGLE17aaj have been identified as the first members of the class based on their unique optical spectroscopic features. Later, the peculiar X-ray transient AT2019avd (Trakhtenbrot et al. 2020) was found with remarkable common properties, however, its classification as an AT2017bgt-like transient is unclear. A more promising case is AT2021loi (Makrygianni et al. 2023), a UV-bright transient event with very similar multi-wavelength properties to the instances studied by Trakhtenbrot et al. (2019). These events show unobscured AGN-like optical spectra with significant and persistent BF lines, extremely strong UV emission and slowly, on yearly scales, decaying optical emission. These AGN are characterized by significantly and longer-term intensified accretion. Hereafter, we will refer to this new class of SMBH flares as Bowen Fluorescence Flares (BFFs), following Makrygianni et al. (2023). Using the Zwicky Transient Facility (ZTF, Masci et al. 2019) Public Survey data, Dgany et al. (2023) found only one BFF candidate (AT2021seu) among spectroscopically classified transients, suggesting that these flares are rare. Two other candidates, namely AT2022fpx, (Koljonen et al. 2024) and AT2020afhd (Arcavi et al. 2024) were found later.
@@
In recent years, two nuclear transients were found to be spatially coincident with neutrino events detected by the IceCube Neutrino Observatory based on ZTF follow-up campaigns of IceCube neutrino alerts (Stein et al. 2023). The TDE AT2019dsg is likely associated with the neutrino, IC191001A (Stein et al. 2021) and a further promising association, between the candidate TDE AT2019fdr and IC200530A (Reusch et al. 2022) points to TDEs as a potential new class of sources of the extragalactic neutrinos. Both sources were identified as having luminous mid-infrared emission, and a search for similar flares (van Velzen et al. 2024) led to the identification of the transient event AT2019aalc (a.k.a. ZTF19aaejtoy) with the neutrino alert IC191119A. In addition, two candidate obscured TDEs (Jiang et al. 2023) and the extremely energetic candidate TDE AT2021lwx (Yuan et al. 2024) were associated with high-energy neutrinos. A recently claimed association between a neutrino flare detected by the IceCube Observatory and an X-ray flare from the TDE ATLAS17jrp suggests that the X-ray emission of these sources play a crucial rule in producing neutrinos (Li et al. 2024).
@@
The paper is organized as follows. We explain the discovery of the transient event and give an insight about its host galaxy in Sect. 2, summarize our observing campaign and the data reduction in Sect. 3, present the results of the observations in Sect. 4 and discuss them in Sect. 5 and finally give a summary and conclude in Sect. 6. Throughout the paper, we adopt a flat CDM cosmological model with parameters , , and . In this model (Wright 2006), mas angular size corresponds to pc projected linear size at the source redshift (Ahn et al. 2012). This redshift corresponds to a luminosity distance of Mpc. All magnitudes are given in AB system (Oke 1974).
@@ 2 The transient event: AT2019aalc
An analysis of the infrared properties of AT2019dsg and AT2019fdr revealed a strong reverberation signal i.e. the infrared dust echo, known as reprocessed emission of optical/UV photons in the infrared by hot dust around the SMBH, at a distance of pc (Jiang et al. 2016; Dou et al. 2016; Lu et al. 2016; van Velzen et al. 2021b; Jiang et al. 2021; van Velzen et al. 2024). In addition to the exceptionally high infrared luminosities of these transients, the neutrino detection times appear to be temporally consistent with the infrared luminosity peaks (Stein et al. 2021; Reusch et al. 2022). This discovery motivated a systematic search for neutrino emission using an extended sample of black hole flares (van Velzen et al. 2024). This archival search revealed the nuclear transient AT2019aalc with a powerful optical flare and dust echo emission in spatial coincidence with the high-energy neutrino event, IC191119A (IceCube Collaboration 2019) detected by the IceCube Observatory.
@@
The host galaxy of AT2019aalc (2MASX J15241665+0451192) is a barred spiral (SBa) galaxy hosting an active nucleus (Oh et al. 2013). Based on its Sloan Digital Sky Survey (SDSS: Abazajian et al. 2009) spectrum it was further classified as a broad-line Seyfert-1 galaxy (Trump et al. 2013; Liu et al. 2019). The galaxy was detected in archival radio sky surveys e.g., the Very Large Array Faint Images of the Radio Sky at Twenty-cm (VLA FIRST: Condon et al. 1998). Fig. 1 shows the Pan-STARRS (Chambers et al. 2016) i-band optical image of the galaxy together with the contours of the FIRST radio data (at GHz) and the position of the X-ray counterpart detected by Swift/XRT in . The galaxy is not present in any X-ray catalogs prior to the optical discovery of the transient in .
@@
In addition to the significant IR echo, AT2019aalc shares multi-wavelength characteristics with the two other neutrino candidate transients: a soft X-ray spectrum and transient radio emission (van Velzen et al. 2024). Moreover, similarly to the other two instances, the infrared peak of the transient event is temporarily coincident with the detection time of the likely associated neutrino event. Notably, Winter & Lunardini (2023) estimated the highest neutrino fluence for AT2019aalc among the studied candidates, due to its high estimated SMBH mass (, van Velzen et al. 2024) and low redshift. Since the proposed association with the high-energy neutrino event was announced roughly years after the discovery of the transient event, the source was not at the focus of attention. Thus, the lack of spectroscopic monitoring during the first optical flare hindered a clear classification of the event. Nevertheless, out of the more than AGN detected by the ZTF, less than show similarly rapid and large outbursts (Reusch et al. 2022; van Velzen et al. 2024) suggesting a very unusual transient. Interestingly, roughly years after the first flare, a significant optical re-brightening started in mid-May 2023 (Veres et al. 2023), illustrated by the long-term ZTF difference light curve shown in Fig. 2.
@@ 3.1.1 ZTF
ZTF forced photometry at the location of AT2019aalc was obtained following the recommendations outlined in Masci et al. (2023). We corrected for extinction by adopting a galactic extinction of (Schlafly & Finkbeiner 2011). As shown in Fig. 2, the forced photometry yields a negative flux in the period between the two flares (and also just before the first flare). This implies that during these epochs, the flux decreased below the mean flux of the reference images. For the -band these were obtained between March and May, 2018. While the baseline used to obtain the difference flux of the -band is based on images obtained between March and September, 2018.
@@ 3.1.2 UVOT
The Neil Gehrels Swift Observatory (Arnaud 1996a) performed Target of Opportunity (ToO) observations (project codes: 18972, 19188, 19993, 20997 and 21692, Proposers: Reusch, Veres & Sniegowska) on AT2019aalc from 2023-06-23 to 2025-01-02 with a total exposure time of ks. The Ultraviolet and Optical Telescope (UVOT, Roming et al. 2005) onboard the satellite observed the target field with the following four filters in each epoch: UW2 (central wavelength, 1928 ), UVM2 (2246 ), UW1 (2600 ) and U (3465 ). We analyzed the UVOT images using the Swift UVOT tools included in the heasoft 6.30 software package. We measured the flux using uvotsource, applying a circular aperature of arcseconds (which captures nearly all the flux of the host galaxy). The background was measured using four 6 arcsecond regions in four quadrants away from the target source. To estimate the UV flux before the flare, we use GALEX (Martin et al. 2005) observations obtained in 2007 and 2011. We use the same 15” circular aperture that was applied to the Swift/UVOT observations and use the gPhoton software (Million et al. 2016) to extract the flux. Over the 4 year baseline of the GALEX observations, the NUV magnitudes change by about 0.3 mag. Following van Velzen et al. (2021a), we combine the mean GALEX flux with the SDSS (York et al. 2000) magnitudes and apply Prospector (Johnson & Leja 2017) to find the best-fit stellar population synthesis model (Conroy et al. 2009) that describes these data. From this model we estimate the baseline flux in the UVOT filters and this baseline is subtracted to obtain the difference flux. The baseline flux in the UVM2 band is 17.2 mag, implying an increase of 1.8 mag relative to the peak of the UVM2 light curve. This large increase implies that, at the peak of the flare, we are not sensitive to the value of the baseline flux. This is useful because the baseline of the ZTF forced photometry was obtained around 2018 (at the start of the ZTF survey), while the baseline that is used to obtain the UVOT difference flux is mainly determined by the GALEX observations that were obtained a decade earlier.